# Tractable Algorithms for Changepoint Detection in Player Performance Metrics

### v1

@@ 5.1 Plate discipline results
The MLB regular season ran from March 30 to October 1 in 2023 and March 20 (for a two game series in South Korea between the Padres and Dodgers before the rest of the games began on March 28) to September 30 in 2024. For chase rate and whiff rate respectively, 62% and 64% of the flagged changepoints occured in May through August.
@@ 6 Detecting changes in pitcher fastball velocity
Next we consider the problem of detecting changes in a pitcher’s primary fastball velocity. We assume that the velocity of a pitcher’s $i$th fastball follows a normal distribution, consistent with empirical pitch-level data (MLB Advanced Media 2025): $y_{i}\sim\mathcal{N}(\mu,\sigma^{2})$ for some mean $\mu$ and variance $\sigma^{2}$.
@@
For metrics such as fastball velocity, where within-pitcher variability is low, even small fluctuations can appear statistically significant, especially with large sample sizes. However, these minor changes are often not meaningful from a baseball perspective. By incorporating a location shift $\Delta$ into the null hypothesis (e.g., testing whether the post-change average differs by more than 1 mph), we focus detection on changes that are practically significant. This approach reduces false positives and ensures that only meaningful deviations, such as those indicative of injury, mechanical adjustment, or performance decline, are flagged as changepoints. For low-variance metrics like velocity, where the difference between statistical and practical significance is especially pronounced, this modification improves interpretability and aligns inference with domain expertise.
@@ 1 No 30 (21%)
7 Ground-truth evaluation: pitcher role changes
@@
It is difficult to construct a ground truth changepoint dataset for many of these types of changes because we do not know, for example, that a pitcher’s fastball velocity will definitely decrease as he ages from 32 to 33. We often also do not know if a pitcher has made a biomechanical change (unless, for example, they tell the media or we have inside information). However, the last bullet point provides an avenue for constructing a ground-truth dataset to test our changepoint detection algorithm, as we know when a pitcher experiences a role change (e.g., from starter to reliever). In this section we evaluate our algorithm’s ability to detect known changes in pitcher statistics due to a role change using a curated ground-truth set of pitcher role changes.
We construct a dataset of pitchers, in the 2023 and 2024 MLB seasons, that switched from a relief to starting pitcher role to tune and test our changepoint detection algorithms. We consider fastball velocity, as it is well established that pitchers tend to lose velocity when transitioning from a relief role to a starting role (and gain velocity when moving from starter to reliever), since outings with a higher pitch volume typically require pitchers to reduce their intensity in order to remain effective over a longer duration compared to shorter relief appearances (Choi 2022, Palattella 2023).
@@
We focus on building a groundtruth dataset around reliever to starter transitions as relievers frequently transition to a starting role due to improved performance and stamina, but face a drop in velocity after transitioning due to the endurance required to pitch through longer outings. On the other hand starters transitioning to relievers may switch roles due to a decline in performance and the boost to their velocity that comes from the role change might be washed out due to their performance decline.
To identify players that transition from a relief to starting role, we look at players that entered the game in the 5th inning or later (indicative of relief outings) more than 10 times and pitched more than 4 innings (indicative of a start) more than 10 times across the 2023 and 2024 seasons. We then filtered this list to only consider pitchers that increased their average number of innings pitched from 2023 to 2024 (i.e., moved from a reliever to starter). We also excluded pitchers that transitioned roles more than once during this time period. We confirmed and supplemented this list in consultation with news articles highlighting reliever to starter transitions (Murphy 2024, Adams 2024).
@@
Figure 3: Jordan Hick’s and Jose Soriano’s sinker velocity by pitch for MLB games in 2023 and 2024. Each plot features two horizontal blue lines. The leftmost and rightmost horizontal lines represent the average sinker velocity for each pitcher as a reliever and starter respectively. Hick’s average sinker velocity decreases by 5.5 mph after the transition. Soriano’s average sinker velocity increases by 0.12 mph after the transition.
@@ 8 Discussion
In this paper we present methods for quantifying metric stabilization, as well as tractable algorithms for detecting changepoints in binary and continuous player performance metrics. In particular, we propose the use of a likelihood based changepoint detection algorithm combined with split sample inference, using either nonparametric tests or tests appropriate to the underlying data distribution. Furthermore, hypothesis tests can incorporate a shift parameter, allowing users to only flag changes larger than a pre-specified magnitude. We demonstrate the application of this algorithm on batter plate discipline metrics and pitcher fastball velocity. Our method is able to flag velocity changes in 91% of pitchers that transitioned from a relief to starting role in the 2023 or 2024 MLB season. Furthermore, we find that for some metrics more than 60% of detected changes occur in-season.
@@
While this work draws on existing changepoint detection literature, methods are combined in a way specifically tailored to this setting. In particular, the inclusion of split-sample inference with hypothesis tests that allow for a shift parameter allows users to better control the Type I error rate as well as have more control over the magnitude of changes detected. Furthermore, existing R packages for changepoint detection generally do not handle binary data well, as calculations are typically based on the normal distribution (Killick & Eckley 2014, Lindeløv 2024). We provide an R implementation of all methods at https://github.com/akglazer/sports-changepoint.
We chose to focus on plate discipline metrics and fastball velocity as those metrics are less likely to be influenced by other factors such as the opponent and ballpark. However, fastball velocity can experience small fluctuations throughout a game as a pitcher fatigues. Future research could investigate the impact of fatigue on pitcher performance and integrate this into changepoint detection methods.
@@
In this paper we primarily focus on detecting changes in means, however, the methods could easily be extended to look for changes in variance or other distributional characteristics. Furthermore, a nonparametric approach could be employed instead of the likelihood ratio-based method when the underlying data distribution is unknown. For example, one could use a divergence measure, as in Matteson & James (2014), instead of a likelihood ratio test statistic.
@@
While we focus on split sample inference as a way to reduce the false positive rate, alternative approaches could also be considered. One such method, proposed by Jewell et al. (2022), evaluates the probability, under the null hypothesis of no true change in mean at a candidate changepoint, that the observed difference in means between segments on either side of a candidate changepoint could arise by chance. Specifically, it conditions on the set of changepoints and assesses whether the observed change is unusually large relative to all datasets that would yield the same changepoint configuration.
@@
While we focus on baseball applications in this paper, this work could easily be applied to other areas where it is of interest to detect changes in performance, e.g., manufacturing (Shang et al. 2017), healthcare or finance.

### v2

@@ 3.3 Multiple changepoint detection
4 Simulation studies
In this section, we evaluate the performance of the proposed changepoint detection framework via simulation. The goals of this section are to (i) evaluate whether the split sample procedure adequately controls type I error rate; and (ii) examine the power of the algorithm to detect a changepoint when a true change is present. We consider both binary and continuous data-generating processes, mirroring the plate discipline and fastball velocity applications presented in Sections 5 and 6. Across all simulations, a changepoint is declared only if the full procedure described in Algorithm 1 rejects the null hypothesis at level $\alpha=0.05$ using the held-out sample. Results are based on $1,000$ independent Monte Carlo replicates for each configuration.
4.1 Type I error control
We first evaluate Type I error control under the null hypothesis of no changepoint. A Type I error is recorded if the procedure declares a changepoint at level $\alpha=0.05$ when no change is present. We compare two variants of the pipeline: (i) with sample splitting, as in Algorithm 1, where the scan uses odd-indexed observations and the confirmation test uses even-indexed observations; and (ii) without sample splitting, where the same observations are used both to select the candidate changepoint and to perform the confirmation test.
Continuous null model.
We generate $T$ observations from a stationary Gaussian model with constant mean,
$y_{i}\sim\mathcal{N}(\mu,\sigma^{2}),\quad i=1,\dots,T,$
with $\mu=93$ and $\sigma=1$ and $T\in\{400,800\}$. For each replicate, we compute the likelihood-ratio scan statistic and candidate changepoint using the Gaussian likelihood, and apply the two-sample permutation test at level $\alpha$ to confirm the candidate changepoint. Under sample splitting, the scan is performed on odd-indexed observations and the test on even-indexed observations; without sample splitting, both steps use the full sequence.
Binary null model.
We generate binary data as independent Bernoulli trials,
$y_{i}\sim\text{Bernoulli}(p),\quad i=1,\dots,T,$
with $p\in\{0.10,0.40\}$ and $T\in\{400,800\}$. Candidate changepoints are obtained via the Bernoulli likelihood-ratio scan, and confirmation is performed using Fisher’s Exact Test at level $\alpha$, with and without sample splitting as above.
Table 1 reports empirical rejection rates. Across both data types, sample splitting yields empirical Type I error near the nominal level, whereas omitting sample splitting substantially inflates false positives.
Sensitivity to mild serial dependence.
To assess robustness to short-range dependence, we simulate sequences under two types of dependence.
For continuous metrics, we generate autocorrelated sequences by simulating $AR(1)$ processes:
$y_{t}=\mu+\rho(y_{t-1}-\mu)+\varepsilon_{t},\qquad\varepsilon_{t}\overset{iid}{\sim}\mathcal{N}(0,1),$
with $\mu=93$ and $\rho=0.1$.
For binary sequences, we simulate from a first-order Markov model with constant marginal success probability $p$:
$\Pr(y_{t}=1\mid y_{t-1})=\begin{cases}p+\rho,&y_{t-1}=1,\\ p-\rho,&y_{t-1}=0,\end{cases}$
with $p=0.25$ and $\rho=0.05$. We repeat the Type I error experiments for $T=800$.
Results are reported in Table 1 and show that sample splitting substantially reduces false positives relative to the no-splitting baseline. Type I error is well controlled for binary sequences under mild dependence, while for continuous sequences sample splitting substantially mitigates, but does not fully eliminate, Type I error inflation under serial dependence.
Table 1: Empirical Type I error rates under the null hypothesis of no changepoint, comparing the proposed sample-splitting procedure to a no-splitting baseline.
Metric Setting With splitting Without splitting
Continuous $T=400$ 0.049 0.479
Continuous $T=800$ 0.051 0.533
Continuous $T=800,AR(1)$ ($\rho=0.1$) 0.081 0.623
Binary ($p=0.10$) $T=400$ 0.037 0.452
Binary ($p=0.40$) $T=400$ 0.045 0.486
Binary ($p=0.10$) $T=800$ 0.035 0.478
Binary ($p=0.40$) $T=800$ 0.049 0.545
Binary ($p=0.25$) $T=800$, Markov dep. $(\rho=0.05)$ 0.051 0.674
4.2 Power
Next we evaluate the power of the algorithm to detect a changepoint when a true change is present. Power comparisons are restricted to the sample-splitting procedure, as the no-splitting baseline does not maintain nominal Type I error under the null, as demonstrated in the previous section.
Continuous metrics.
We generate sequences of length $T\in\{400,800\}$ from a Gaussian mean-shift model with a single changepoint at $c=\lfloor T/2\rfloor$:
$y_{t}\sim\begin{cases}\mathcal{N}(\mu,\sigma^{2}),&t\leq c,\\ \mathcal{N}(\mu+\delta,\sigma^{2}),&t>c,\end{cases}$
with $\mu=93$ and $\sigma=1$. We consider effect sizes $\delta\in\{0.25,0.5\}$. The confirmation step uses a two-sample permutation test.
Table 2 shows that power increases rapidly with both the magnitude of the mean shift and the length of the observed sequence. For small effect sizes ($\delta=0.25$), power is limited for shorter sequences but improves substantially as $T$ increases, reflecting the reduced variance in the split-sample confirmation step. For larger, practically meaningful shifts ($\delta=0.5$), the procedure achieves high power even at $T=400$ and is nearly certain to detect the change by $T=800$.
Table 2: Empirical power for the continuous mean-shift model.
$T$ $\delta$ Power
400 0.25 0.318
400 0.5 0.88
800 0.25 0.549
800 0.5 0.988
Binary metrics.
Binary sequences of length $T\in\{800,1600\}$ are generated from a Bernoulli rate-shift model with a changepoint at $c=\lfloor T/2\rfloor$:
$y_{t}\sim\begin{cases}\text{Bernoulli}(p_{1}),&t\leq c,\\ \text{Bernoulli}(p_{2}),&t>c,\end{cases}$
with $p_{1}=0.20-\delta$ and $p_{2}=0.20+\delta$ for $\delta\in\{0.03,0.05,0.08\}$. Candidate changepoints are obtained via the Bernoulli likelihood-ratio scan, and confirmation uses Fisher’s Exact Test at level $\alpha=0.05$. Results are summarized in Table 3.
As expected, power depends strongly on both the magnitude of the rate shift and the sequence length. Small shifts ($\delta=0.03$) are difficult to detect, particularly for shorter sequences, reflecting the inherent noisiness of binary outcomes and the reduced effective sample size induced by split-sample inference. However, power increases rapidly with both $T$ and $\delta$: for moderate shifts ($\delta=0.05$), power exceeds 80% by $T=1600$, and for larger, practically meaningful shifts ($\delta=0.08$), detection is very high even at $T=800$.
The reduced sensitivity for small shifts is an explicit trade-off of the split-sample design, which prioritizes Type I error control over aggressive detection. In applied MLB settings, this trade-off is appropriate: typical chase-rate sequences exceed 1,000 pitches per season after filtering, and changes of interest are often larger than a few percentage points. Thus, the $T=1600$ results are representative of practical use, where the method exhibits strong power while maintaining reliable false positive control.
Table 3: Empirical power for the binary rate-shift model.
$T$ $\delta$ Power
800 0.03 0.151
800 0.05 0.479
800 0.08 0.916
1600 0.03 0.309
1600 0.05 0.811
1600 0.08 0.996
@@ 5.1 Plate discipline results
The MLB regular season ran from March 30 to October 1 in 2023 and March 20 (for a two game series in South Korea between the Padres and Dodgers before the rest of the games began on March 28) to September 30 in 2024. For chase rate and whiff rate respectively, 62% and 64% of the flagged changepoints occurred in May through August.
@@ 6 Detecting changes in pitcher fastball velocity
Next we consider the problem of detecting changes in a pitcher’s primary fastball velocity. We assume that the velocity of a pitcher’s $i$th fastball follows a normal distribution, consistent with empirical pitch-level data (statcastVeloDist): $y_{i}\sim\mathcal{N}(\mu,\sigma^{2})$ for some mean $\mu$ and variance $\sigma^{2}$.
@@
For metrics such as fastball velocity, where within-pitcher variability is low, even small changes can be detected with high statistical power when sufficient data are available, as demonstrated in our simulation studies in Section 4. However, small fluctuations are often not meaningful from a baseball perspective, especially with large sample sizes where statistical significance can outpace practical relevance. By incorporating a location shift $\Delta$ into the null hypothesis (e.g., testing whether the post-change average differs by more than 1 mph), we focus detection on changes that are practically significant. This modification reduces false positives driven by trivial variation and ensures that flagged changepoints correspond to deviations likely to reflect substantive changes, such as injury, mechanical adjustment, or sustained performance decline. For low-variance metrics like velocity, where the distinction between statistical and practical significance is especially pronounced, this approach improves interpretability and better aligns inference with domain expertise.
@@ 1 No 30 (21%)
7 Validation using known pitcher role transitions
@@
It is difficult to construct a ground truth changepoint dataset for many of these types of changes because we do not know, for example, that a pitcher’s fastball velocity will definitely decrease as he ages from 32 to 33. We often also do not know if a pitcher has made a biomechanical change (unless, for example, they tell the media or we have inside information). However, the last bullet point provides an avenue for constructing a quasi-ground-truth dataset to test our changepoint detection algorithm, as we know when a pitcher experiences a role change (e.g., from starter to reliever). In this section we evaluate our algorithm’s ability to detect known changes in pitcher statistics due to a role change using a curated ground-truth set of pitcher role changes.
We construct a dataset of pitchers, in the 2023 and 2024 MLB seasons, that switched from a relief to starting pitcher role to tune and test our changepoint detection algorithms. We consider fastball velocity, as it is well established that pitchers tend to lose velocity when transitioning from a relief role to a starting role (and gain velocity when moving from starter to reliever), since outings with a higher pitch volume typically require pitchers to reduce their intensity in order to remain effective over a longer duration compared to shorter relief appearances (choiStarterReliever, palattella23).
@@
We focus on building a quasi-groundtruth dataset around reliever to starter transitions as relievers frequently transition to a starting role due to improved performance and stamina, but face a drop in velocity after transitioning due to the endurance required to pitch through longer outings. On the other hand starters transitioning to relievers may switch roles due to a decline in performance and the boost to their velocity that comes from the role change might be washed out due to their performance decline.
To identify players that transition from a relief to starting role, we look at players that entered the game in the 5th inning or later (indicative of relief outings) more than 10 times and pitched more than 4 innings (indicative of a start) more than 10 times across the 2023 and 2024 seasons. We then filtered this list to only consider pitchers that increased their average number of innings pitched from 2023 to 2024 (i.e., moved from a reliever to starter). We also excluded pitchers that transitioned roles more than once during this time period. We confirmed and supplemented this list in consultation with news articles highlighting reliever to starter transitions (relieverToStarter, relieverToStarter2).
@@
Figure 2: Jordan Hicks’ and Jose Soriano’s sinker velocity by pitch for MLB games in 2023 and 2024. Each plot features two horizontal blue lines. The leftmost and rightmost horizontal lines represent the average sinker velocity for each pitcher as a reliever and starter respectively. Hicks’ average sinker velocity decreases by 5.5 mph after the transition. Soriano’s average sinker velocity increases by 0.12 mph after the transition.
@@ 8 Discussion
In this paper we present tractable methods for detecting changepoints in binary and continuous player performance metrics. In particular, we propose a likelihood-based changepoint detection framework combined with split-sample inference, using either nonparametric tests or tests appropriate to the underlying data distribution. Hypothesis tests can additionally incorporate a shift parameter, allowing users to restrict attention to changes exceeding a pre-specified magnitude.
Through simulation studies, we show that the proposed procedure achieves reliable Type I error control under a range of data-generating settings and retains meaningful power to detect practically relevant changes. We then demonstrate the utility of the method on MLB data, applying it to batter plate discipline metrics and pitcher fastball velocity. Our approach flags velocity changes in 91% of pitchers who transitioned from a relief to starting role during the 2023 or 2024 MLB seasons. Furthermore, we find that for some metrics more than 60% of detected changes occur in-season.
@@
While this work draws on existing changepoint detection literature, methods are combined in a way specifically tailored to this setting. In particular, the inclusion of split-sample inference with hypothesis tests that allow for a shift parameter allows users to better control the Type I error rate as well as have more control over the magnitude of changes detected. Furthermore, we expand on existing R packages for changepoint detection (killick2014changepoint, lindeloev2024mcp_packages) by providing functionality for binary sequences and split sample inference. We provide an R implementation of all methods at https://github.com/akglazer/sports-changepoint.
We chose to focus on plate discipline metrics and fastball velocity because these metrics are commonly monitored by teams and are less directly influenced by contextual aggregation than outcome-based statistics such as batting average or earned run average. Nonetheless, these metrics are not fully context-free. For example, fastball velocity can exhibit some within-game fluctuations due to fatigue or environmental conditions. While our simulation studies suggest that the proposed split-sample procedure maintains Type I error control under mild dependence, future work could more explicitly model temporal dependence and contextual factors within the changepoint detection framework.
@@
In this paper we primarily focus on detecting changes in means, however, the methods could easily be extended to look for changes in variance or other distributional characteristics. Furthermore, a nonparametric approach could be employed instead of the likelihood ratio-based method when the underlying data distribution is unknown. For example, one could use a divergence measure, as in matteson2014nonparametric, instead of a likelihood ratio test statistic.
While permutation tests provide a flexible, distribution-free approach for confirming changepoints, they can be computationally intensive when applied repeatedly across many players and metrics. In applications where distributional assumptions are reasonable, the confirmation step could instead use a parametric test, substantially reducing runtime while preserving the split-sample inference framework.
@@
Although we focus on split sample inference as a way to reduce the false positive rate, alternative approaches could also be considered. One such method, proposed by jewell2022testing, evaluates the probability, under the null hypothesis of no true change in mean at a candidate changepoint, that the observed difference in means between segments on either side of a candidate changepoint could arise by chance. Specifically, it conditions on the set of changepoints and assesses whether the observed change is unusually large relative to all datasets that would yield the same changepoint configuration.
@@
While we focus on baseball applications in this paper, this work could easily be applied to other areas where it is of interest to detect changes in performance, e.g., manufacturing, healthcare or finance (reeves2007review, shang2017nonparametric, banerjee2020change, kim2022unsupervised).